# Next-to-leading-order electroweak correction to $H\to Z^{0}\gamma$

### v1

@@ Abstract
Inspired by the recent observation of the Higgs boson radiative decay into $Z^{0}$ by ATLAS and CMS Collaborations, we investigate the next-to-leading-order (NLO) electroweak correction to this rare decay process in Standard Model (SM). Implementing the on-shell renormalization scheme, we find that the magnitude of the NLO electroweak correction may reach $7\%$ of the leading order (LO) prediction, much more significant than that of the NLO QCD correction, which is merely about $0.3\%$. After incorporating the ${\cal O}(\alpha)$ correction, the predicted partial width from various $\alpha$ schemes tend to converge to each other. Including both NLO electroweak and QCD corrections, we present the most accurate SM prediction for the branching fraction of this rare decay channel to be ${\cal B}_{\rm th}[H\to Z^{0}\gamma]=(1.55\pm 0.06)\times 10^{-3}$, considerably lower than the measured value ${\cal B}_{\rm exp}[H\to Z^{0}\gamma]=(3.4\pm 1.1)\times 10^{-3}$. Resolving this alarming discrepancy clearly calls for further theoretical investigations, and, more importantly, experimental efforts from HL-LHC and the prospective Higgs factories such as CEPC and FCC.
@@
Recently, the ATLAS and CMS Collaborations jointly reported the evidence of observing the rare decay $H\to Z^{0}\gamma$ ATLAS:2023yqk ; ATLAS:2020qcv ; CMS:2022ahq , and found $\mu_{Z\gamma}=2.2\pm 0.7$, with $\mu\equiv(\sigma{\cal B})_{\rm meas}/(\sigma{\cal B})_{\rm th}$ being a parameter to gauge how well the measurement conforms to the SM prediction. The measured branching fraction is ${\cal B}(H\to Z\gamma)=(3.4\pm 1.1)\times 10^{-3}$, more than twice larger than the SM prediction $(1.5\pm 0.1)\times 10^{-3}$ Djouadi:1997yw ; LHCHiggsCrossSectionWorkingGroup:2016ypw . This should be contrasted with a similar channel $H\to\gamma\gamma$ with $\mu_{\gamma\gamma}=1.10\pm 0.07$, which exhibits a satisfactory agreement between the measurement and the SM prediction. Recently there has appeared a flurry of work attempting to alleviate the discrepancy between experiment and theory for $H\to Z^{0}\gamma$, by invoking various BSM models Hong:2023mwr ; Panghal:2023iqd ; Buccioni:2023qnt ; Boto:2023bpg ; Cheung:2024kml ; He:2024sma .
@@
Needless to say, in order to ultimately resolve the discrepancy related to $H\to Z^{0}\gamma$, it is compulsive to first have a as precise as possible prediction in SM. At leading order (LO), the process $H\to Z^{0}\gamma$ is mediated via a heavy quark or a $W$ boson loop, as depicted in Fig. 1. The LO prediction has been available four decades ago Cahn:1978nz ; Bergstrom:1985hp . It turns out that the $W$-loop induced contribution dominates the $t$-loop induced one, while there exists a destructive interference between these two channels. The NLO QCD correction in numerical form has been known in early 90s Spira:1991tj , and its analytical form has also been available about one decade ago Gehrmann:2015dua ; Bonciani:2015eua . The QCD correction appears to be rather insignificant. This may be partly be understood by the fact that the gluons can only feel the top quark loop, which by itself only yields a small portion of the contribution with respect to the $W$ loop.
Therefore, the missing NLO electroweak correction is envisaged to constitute the major theoretical uncertainty for this decay channel, which has been estimated to reach $5\%$ level of the LO contribution LHCHiggsCrossSectionWorkingGroup:2016ypw . It is the goal of this work to conduct a complete investigation of the NLO electroweak correction to $H\to Z^{0}\gamma$.
@@
where $p_{1}$ and $p_{2}$ signify the momenta of the outgoing $Z$ boson and photon, respectively, and $T_{i}$ ($i=1,\cdots 6$) signify six scalar form factors. The transversity condition $p_{1}\cdot\epsilon_{Z}=p_{2}\cdot\epsilon_{\gamma}=0$ implies that the terms entailing $T_{2}$ and $T_{3}$ do not contribute. Ward identity $p_{2\nu}T^{\mu\nu}=0$ implies that $T_{1}=0$ and $T_{4}=-T_{5}$. $T_{6}$ first starts at two-loop electroweak correction, which can thus be safely neglected as far as we concentrate on the NLO electroweak correction. Therefore only one independent form factor, $T_{5}=-T_{4}$, survives. In practice, it is convenient to introduce a new dimensionless form factor $\widetilde{T}_{5}$:
@@
where the three-momentum of the $Z^{0}$ boson in the Higgs boson rest frame is $|{\bf p}_{1}|={(M_{H}^{2}-M_{Z}^{2})}/{2M_{H}}$. Our main task is to deduce $\widetilde{T}_{5}$ through NLO in $\alpha$.
Outline of calculation. Throughout this work we adopt the Feynman gauge and employ the dimensional regularization to regularize the potential UV divergence. We utilize the package FeynArts Hahn:2000kx to generate the Feynman diagrams and the corresponding amplitudes for $H\to Z^{0}\gamma$ through NLO in $\alpha$. This process entails approximately 50 diagrams at one-loop level, and more than $10^{4}$ diagrams at two-loop level. Some representative diagrams relevant to the NLO QCD and electroweak corrections are depicted in Fig. 2. The form factors $T_{4}$ and $T_{5}$ are extracted from the amplitude with the aid of the covariant projectors Bonciani:2015eua . We employ the packages FeynCalc Mertig:1990an and FormLink Feng:2012tk to perform Lorentz contraction and Dirac matrix trace. We use the packages Apart Feng:2012iq and FIRE Smirnov:2014hma for integration-by-part reduction. We end up with 8 one-loop master integrals (MIs) and over 700 two-loop MIs. The package AMFlow Liu:2017jxz ; Liu:2022mfb ; Liu:2022chg is invoked to evaluate all these MIs with high numerical accuracy.
Figure 2: Some representative two-loop diagrams for $H\to Z^{0}\gamma$. $a)$ represents a sample diagram for NLO QCD correction, while $b)$-$h)$ represent sample diagrams for the NLO electroweak correction.
The on-shell renormalization scheme has been widely used in the field of electroweak radiative correction to tame the UV divergence Ross:1973fp ; Hollik:1988ii . The renormalized parameters are chosen to be those measured very precisely, such as Higgs boson mass, $W/Z$ masses, top quark mass, and QED coupling at Thomson limit. Having implemented the mass and charge renormalization ($e_{0}=Z_{e}e$), we find it convenient to stay with the bare field in the practical calculation. Following the LSZ reduction formula, we multiply the amputated amplitude of $H\to Z^{0}\gamma$ by the factor $\sqrt{Z_{H}Z_{ZZ}Z_{\gamma\gamma}}$, where the $Z_{H}$, $Z_{ZZ}$ and $Z_{\gamma\gamma}$ denote the field-strength renormalization constants associated with $H$, $Z^{0}$ and $\gamma$, respectively. These field strength renormalization constants can be inferred from the unrenormalized propagators to one loop accuracy, by identifying the residues in the on-shell limit.
@@
Three different recipes about charge renormalization. Implementing the charge renormalization in electroweak theory is not unique, which leads to several popular variants in the on-shell renormalization scheme. In the so-called $\alpha(0)$ scheme, where the $\alpha$ is taking the Thomson-limit value, To one-loop order, $\delta Z_{e}$ is expressed as
$\delta Z_{e}|_{\alpha(0)}={1\over 2}{\Pi^{\gamma\gamma}(0)}-{s_{W}\over c_{W}}{\Sigma^{\gamma Z}_{T}(0)\over M_{Z}^{2}},$ (4)
where ${\Pi(s)}\equiv\Sigma_{T}^{\gamma\gamma}(s)/s$. The photon vacuum polarization is sensitive to the low-energy hadronic contribution, thereby an intrinsically non-perturbative quantity. Alternatively, one may rewrite $\delta{Z_{e}}$ in $\alpha(0)$ scheme as
$\delta{Z_{e}}\big{|}_{\alpha(0)}$ $=$ ${1\over 2}\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})+{1\over 2}{\rm Re}\,\Pi^{\gamma\gamma(5)}(M_{Z}^{2})$ (5)
$+$ ${1\over 2}\Pi^{\gamma\gamma}_{\rm rem}(0)-{s_{W}\over c_{W}}{\Sigma^{\gamma Z}_{T}(0)\over M_{Z}^{2}},$
where $\Pi^{\gamma\gamma(5)}(M_{Z}^{2})$ is the photon vacuum polarization from five massless quarks at momentum transfer $M_{Z}^{2}$. $\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})$, being a non-perturbative parameter, absorbs the hadronic contribution to photon vacuum polarization, which can be determined from the measured $R$ values in low-energy $e^{+}e^{-}$ experiments. $\Pi^{\gamma\gamma}_{\rm rem}(0)$ represents the vacuum polarization from the $W$ boson, charged leptons and top quark at zero momentum transfer. Note all these terms except $\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})$ can be computed reliably in perturbation theory.
Two other popular variants of the on-shell renormalization scheme are the $\alpha(M_{Z})$ and $G_{\mu}$ schemes. The corresponding charge renormalization constant can be converted from (4) by $\delta{Z_{e}}\big{|}_{\alpha(M_{Z})}=\delta{Z_{e}}\big{|}_{\alpha(0)}-{1\over 2}\Delta\alpha(M_{Z})$, and $\delta Z_{e}|_{G_{\mu}}=\delta Z_{e}|_{\alpha(0)}-{1\over 2}\Delta r$, respectively. $\Delta\alpha(M_{Z})=\Pi^{\gamma\gamma}_{f\neq t}(0)-{\rm Re}\,\Pi^{\gamma\gamma}_{f\neq t}(M_{Z}^{2})$, and the expression for the oblique parameter $\Delta r$ can be found in Denner:1991kt . The QED coupling constant in the $\alpha(M_{Z})$ and $G_{\mu}$ schemes read
@@
In contrast to the $\alpha(0)$ scheme, these two sub-schemes effectively resum some universal large (non-)logarithmic terms arising from the light fermions and top quark loop.
Numerical predictions. To make concrete predictions, we adopt the following values for the masses of various particles Workman:2022ynf :
$M_{H}=125.25\>{\rm GeV},\quad M_{Z}=91.1876\>{\rm GeV},$
$M_{W}=80.377\>{\rm GeV},\quad m_{t}=172.69\>{\rm GeV},$
$m_{e}=0.5109989\>{\rm MeV},\quad m_{\mu}=105.65837\>{\rm MeV},$
$m_{\tau}=1.77686\>{\rm GeV}.$ (7)
The masses of the charged leptons are retained only for the purpose of computing $Z_{e}$. In all other situations, we only keep the top quark massive and set the remaining five quarks and all charged leptons to be massless.
We adopt the following values of couplings: $\alpha(0)=1/137.035999$, $G_{\mu}=1.1663787\times 10^{-5}~{}{\rm GeV}^{-2}$ and $\Delta\alpha_{\rm had}^{(5)}(M_{Z})=0.02764$ Workman:2022ynf . We also take $\alpha(M_{Z})=1/128.877$ from Refs. Eidelman:1995ny ; Steinhauser:1998rq . With the aid of (6b) one obtains $\alpha_{G_{\mu}}=1/132.168$.
The current measured full width of the Higgs is subject to large uncertainty, $\Gamma_{H}=3.2^{+2.4}_{-1.7}$ MeV Workman:2022ynf . When predicting the branching fraction, we choose to use the much more precise prediction. Thus, it is more convenient to use the theoretically predicted value, $\Gamma_{H}=4.07^{+4.0\%}_{-3.9\%}$ MeV Workman:2022ynf ; LHCHiggsCrossSectionWorkingGroup:2016ypw .
@@
$\alpha(0)$ scheme $\alpha(0)^{3}$ $5.920$ $0.439$ $0.018$ $6.378$ $1.57\pm 0.06$
$\alpha(M_{Z})$ scheme $\alpha(0)\alpha(M_{Z})^{2}$ $6.694$ $-0.465$ $0.021$ $6.249$ $1.54\pm 0.06$
$G_{\mu}$ scheme $\alpha(0)\alpha_{G_{\mu}}^{2}$ $6.364$ $-0.048$ $0.020$ $6.336$ $1.56\pm 0.06$
Democratic scheme $\alpha(0)\alpha(M_{Z})\alpha_{G_{\mu}}$ $6.528$ $-0.249$ $0.020$ $6.299$ $1.55\pm 0.06$
@@
Table 1: Predicted values of the partial width (in units of keV) for $H\to Z^{0}\gamma$ in various $\alpha$ schemes, at various levels of perturbative accuracy. $\Gamma^{\rm LO}$ represents the LO prediction to the decay width. The contributions from the NLO electroweak and QCD corrections are denoted by $\Gamma^{\mathcal{O}(\alpha)}$ and $\Gamma^{\mathcal{O}(\alpha_{s})}$, respectively. We have taken the strong coupling constant $\alpha_{s}(M_{H})=0.113$. $\Gamma^{\rm Sum}$ is obtained by summing the LO contribution together with $\mathcal{O}(\alpha)$ and $\mathcal{O}(\alpha_{s})$ corrections, and the rightmost column is the branching fraction obtained via dividing $\Gamma^{\rm Sum}$ by the full Higgs width, $\Gamma_{H}=4.07^{+4.0\%}_{-3.9\%}$ MeV.
@@
In Appendix A we present the numerical results of the form factor $\widetilde{T}_{5}$ at NLO accuracy in different $\alpha$ schemes. Plugging into (3), we are able to present the predictions for the partial width in various $\alpha$ schemes. It may look natural to take the QED coupling constant at the vertex attached to the outgoing photon as $\alpha(0)$. For the electromagnetic coupling constant that appears elsewhere, one has some freedom to choose different $\alpha$ scheme. In Table 1 we enumerate our predictions for $\Gamma(H\to Z^{0}\gamma)$ (in units of keV) based on four different $\alpha$ schemes at various levels of perturbative accuracy.
@@
As can be seen from Table 1, we confirm that the NLO QCD correction is positive but tiny, which constitutes only $0.3\%$ of the LO partial width Bonciani:2015eua . We also notice that, the NLO electroweak corrections is positive in the $\alpha(0)$ scheme, while negative in other three schemes. In addition, the ${\cal O}(\alpha)$ corrections appear to be sizable in magnitude in both the $\alpha(0)$ and $\alpha(M_{Z})$ schemes, which may approach approximately $7\%$ of the LO prediction. In contrast, the $G_{\mu}$ scheme as well as the Democratic scheme yield a relatively modest ${\cal O}(\alpha)$ correction. Note the LO predictions of the partial width from four different $\alpha$ scheme are scattered in a wide range, from $5.920$ keV to $6.694$ keV. Interestingly, after including the NLO electroweak correction, the scheme dependence becomes substantially reduced, with the relative uncertainties among different schemes less than $2\%$. Taking the predictions from various $\alpha$ schemes as an estimation of the theoretical uncertainty, we then obtain the most accurate SM prediction to be ${\cal B}[H\to Z^{0}\gamma]=(1.55\pm 0.06)\times 10^{-3}$. Note the error in the branching fraction mainly stem from the uncertainty associated with the full width the of the Higgs boson  [^5] We can estimate other potential sources of uncertainty. The contributions from the top quark loop constitute approximately $-10\%$ of the LO decay width. Taking into account the Yukawa coupling strength of $Hb\bar{b}$ is suppressed with respect to $Ht\bar{t}$ by a factor of $m_{b}/m_{t}$, and we estimate that retaining the bottom quark mass introduces a relative error of several per mille. Furthermore, uncertainties in the mass of the top quark and the $W$ boson may also introduce uncertainty of several per mille. All in all, these sorts of uncertainties are of the same order of magnitude as the NLO QCD correction, which are significantly smaller than the uncertainty stemming from four different $\alpha$ schemes..
@@
It is evident that the most accurate SM prediction for the branching fraction is significantly lower than the measured value, ${\cal B}_{\rm exp}[H\to Z^{0}\gamma]=(3.4\pm 1.1)\times 10^{-3}$ ATLAS:2023yqk . It may be too early to claim that some sort of the BSM physics must be invoked to resolve this discrepancy. More accurate measurements from HL-LHC, and the prospective Higgs factories such as CEPC and FCC, will play a crucial role to clear the smoke.

### v3

@@ Abstract
Inspired by the recent observation of the Higgs boson radiative decay into $Z^{0}$ by ATLAS and CMS Collaborations, we investigate the next-to-leading-order (NLO) electroweak correction to this rare decay process in Standard Model (SM). Implementing the on-shell renormalization scheme, we find that the magnitude of the NLO electroweak correction may reach $7\%$ of the leading order (LO) prediction, much more significant than that of the NLO QCD correction, which is merely about $0.3\%$. After incorporating the ${\cal O}(\alpha)$ correction, the predicted partial width from various $\alpha$ schemes tend to converge to each other. Including both NLO electroweak and QCD corrections, the SM prediction for the branching fraction shifts from the LO value of $(1.40-1.71)\times 10^{-3}$ to $(1.55\pm 0.06)\times 10^{-3}$, considerably lower than the measured value ${\cal B}_{\rm exp}[H\to Z^{0}\gamma]=(3.4\pm 1.1)\times 10^{-3}$. Resolving this discrepancy clearly calls for further theoretical investigations, and, more importantly, experimental efforts from HL-LHC and the prospective Higgs factories such as CEPC and FCC-ee.
@@
Recently, the ATLAS and CMS Collaborations jointly reported the evidence of observing the rare decay $H\to Z^{0}\gamma$ ATLAS:2023yqk ; ATLAS:2020qcv ; CMS:2022ahq , and found $\mu_{Z\gamma}=2.2\pm 0.7$, with $\mu\equiv(\sigma{\cal B})_{\rm meas}/(\sigma{\cal B})_{\rm th}$ being a parameter to gauge how well the measurement conforms to the SM prediction. The measured branching fraction is ${\cal B}(H\to Z\gamma)=(3.4\pm 1.1)\times 10^{-3}$, more than twice larger than the SM prediction $(1.5\pm 0.1)\times 10^{-3}$ Djouadi:1997yw ; LHCHiggsCrossSectionWorkingGroup:2016ypw . This should be contrasted with a similar channel $H\to\gamma\gamma$ with $\mu_{\gamma\gamma}=1.04^{+0.10}_{-0.09}$ ATLAS:2022tnm , which exhibits a satisfactory agreement between the measurement and the SM prediction. Recently there has appeared a flurry of theoretical work attempting to alleviate the discrepancy between the measurements and SM predictions for $H\to Z^{0}\gamma$, by invoking various BSM models Herrero:2020dtv ; Hong:2023mwr ; Panghal:2023iqd ; Buccioni:2023qnt ; Barducci:2023zml ; Boto:2023bpg ; Cheung:2024kml ; He:2024sma .
@@
Needless to say, in order to ultimately resolve the discrepancy related to $H\to Z^{0}\gamma$, it is compulsive to first have a as precise as possible prediction in SM. At leading order (LO), the process $H\to Z^{0}\gamma$ is mediated via a heavy quark or a $W$ boson loop, as depicted in Fig. 1. The LO prediction has been available four decades ago Cahn:1978nz ; Bergstrom:1985hp . It turns out that the $W$-loop induced contribution dominates the $t$-loop induced one, while there exists a destructive interference between these two channels. The NLO QCD correction in numerical form has been known in early 90s Spira:1991tj , and its analytical form has also been available about one decade ago Gehrmann:2015dua ; Bonciani:2015eua . The QCD correction appears to be rather insignificant. This may be partly understood by the fact that the gluons can be attached only to the top quark loop, which by itself only yields a small portion of the contribution with respect to the $W$ loop.
Therefore, the missing NLO electroweak correction is envisaged to constitute the major theoretical uncertainty for this decay channel, which has been estimated to reach $5\%$ level of the LO contribution LHCHiggsCrossSectionWorkingGroup:2016ypw . It is the goal of this work to conduct a detailed investigation on the NLO electroweak correction for $H\to Z^{0}\gamma$.
@@
where $p_{1}$ and $p_{2}$ signify the momenta of the outgoing $Z$ boson and photon, respectively, and $T_{i}$ ($i=1,\cdots 6$) signify six scalar form factors. The transversity condition $p_{1}\cdot\epsilon_{Z}=p_{2}\cdot\epsilon_{\gamma}=0$ implies that the terms entailing $T_{2}$ and $T_{3}$ do not contribute. Ward identity $p_{2\nu}T^{\mu\nu}=0$ implies that $T_{1}=0$ and $T_{4}=-T_{5}$. The $CP$-violating $T_{6}$ term first arises at two-loop order in SM, but its interference with the LO amplitude yields a vanishing contribution. Hence we can neglect the $T_{6}$ term at the outset as far as the NLO electroweak correction is concerned. Consequently, only one independent form factor, $T_{5}=-T_{4}$, finally survives. In practice, it is convenient to introduce a new dimensionless form factor $\widetilde{T}_{5}$:
@@
where the three-momentum of the $Z^{0}$ boson in the Higgs boson rest frame is $|{\bf p}_{1}|={(M_{H}^{2}-M_{Z}^{2})}/{2M_{H}}$. Our task is to deduce $\widetilde{T}_{5}$ through NLO in $\alpha$.
Outline of calculation. Throughout this work we adopt the Feynman gauge and employ the dimensional regularization to regularize UV divergences. We utilize the package FeynArts Hahn:2000kx to generate Feynman diagrams and the corresponding amplitudes for $H\to Z^{0}\gamma$ through NLO in $\alpha$. This process entails approximately 50 LO (one-loop) diagrams, and more than $10^{4}$ NLO (two-loop) diagrams (including the NLO QCD diagrams). Some representative diagrams relevant to the NLO QCD and electroweak corrections are depicted in Fig. 2. The form factors $T_{4}$ and $T_{5}$ are extracted from the amplitude with the aid of the covariant projectors Bonciani:2015eua . We employ the packages FeynCalc Mertig:1990an and FormLink Feng:2012tk to perform Lorentz contraction and Dirac trace. We use the packages Apart Feng:2012iq and FIRE Smirnov:2014hma for integration-by-part reduction. We end up with 8 one-loop master integrals (MIs) and over 700 two-loop MIs. The package AMFlow Liu:2017jxz ; Liu:2022mfb ; Liu:2022chg is invoked to evaluate all these MIs with high numerical accuracy.
Figure 2: Some representative two-loop diagrams for $H\to Z^{0}\gamma$. $a)$ represents a sample diagram for NLO QCD correction, while $b)$-$h)$ represent sample diagrams for NLO electroweak correction.
@@
Three different recipes about charge renormalization. Implementing the charge renormalization in electroweak theory is not unique, which leads to several popular variants in the on-shell renormalization scheme. In the so-called $\alpha(0)$ scheme, where $\alpha$ is taking with the Thomson-limit value. To one-loop order, $\delta Z_{e}$ is expressed as
$\delta Z_{e}|_{\alpha(0)}={1\over 2}\Pi^{AA}(0)-{s_{W}\over c_{W}}{\Sigma^{AZ}_{T}(0)\over M_{Z}^{2}},$ (4)
where ${\Pi(s)}\equiv\Sigma_{T}^{AA}(s)/s$. The photon vacuum polarization in small momentum transfer is sensitive to the low-energy hadronic contribution, thereby an intrinsically nonperturbative quantity. Alternatively, one may rewrite $\delta{Z_{e}}$ in $\alpha(0)$ scheme as
$\delta{Z_{e}}\big{|}_{\alpha(0)}$ $=$ ${1\over 2}\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})+{1\over 2}{\rm Re}\,\Pi^{{AA}(5)}(M_{Z}^{2})$ (5)
$+$ ${1\over 2}\Pi^{AA}_{\rm rem}(0)-{s_{W}\over c_{W}}{\Sigma^{AZ}_{T}(0)\over M_{Z}^{2}},$
where $\Pi^{{AA}(5)}(M_{Z}^{2})$ is the photon vacuum polarization from five massless quarks at momentum transfer $M_{Z}^{2}$. $\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})$, as a nonperturbative parameter that encapsulate the hadronic contribution to photon vacuum polarization, can be determined from the measured $R$ values in low-energy $e^{+}e^{-}$ experiments. $\Pi^{AA}_{\rm rem}(0)$ represents the vacuum polarization from the $W$ boson, charged leptons and top quark at zero momentum transfer. Note all these terms except $\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})$ can be computed reliably in perturbation theory.
Two other popular variants of the on-shell renormalization scheme are the $\alpha(M_{Z})$ and $G_{\mu}$ schemes. The corresponding charge renormalization constant can be adapted from (4) into $\delta{Z_{e}}\big{|}_{\alpha(M_{Z})}=\delta{Z_{e}}\big{|}_{\alpha(0)}-{1\over 2}\Delta\alpha(M_{Z})$, and $\delta Z_{e}|_{G_{\mu}}=\delta Z_{e}|_{\alpha(0)}-{1\over 2}\Delta r$, respectively. $\Delta\alpha(M_{Z})=\Pi^{AA}_{f\neq t}(0)-{\rm Re}\,\Pi^{AA}_{f\neq t}(M_{Z}^{2})$, and the expression for the oblique parameter $\Delta r$ can be found in Denner:1991kt . For latter convenience, we split the $\Delta\alpha(M_{Z})$ into two different sources of contributions:
$\Delta\alpha\left(M_{Z}\right)=\Delta\alpha_{q}\left(M_{Z}\right)+\Delta\alpha_{\ell}\left(M_{Z}\right),$ (6)
where $\Delta\alpha_{q}$ and $\Delta\alpha_{\ell}$ represent the contributions from the five light quarks and three charged leptons, respectively. They are defined by
$\Delta\alpha_{q}(M_{Z})$ $=$ $\Pi^{AA}_{f=u,d,s,c,b}(0)-{\rm Re}\,\Pi^{AA}_{f=u,d,s,c,b}(M_{Z}),\phantom{xxx}$ (7a)
$\Delta\alpha_{\ell}(M_{Z})$ $=$ $\Pi^{AA}_{f=e,\mu,\tau}(0)-{\rm Re}\,\Pi^{AA}_{f=e,\mu,\tau}(M_{Z}).$ (7b)
The QED coupling constants in the $\alpha(M_{Z})$ and $G_{\mu}$ schemes become
@@
In contrast to the $\alpha(0)$ scheme, these two sub-schemes effectively resum some universal large (non-)logarithmic terms arising from the light fermion loop and top quark loop.
Numerical predictions. To make concrete predictions, we adopt the following values for the masses of various particles Workman:2022ynf : $M_{H}=125.25\>{\rm GeV}$, $M_{Z}=91.1876\>{\rm GeV}$, $M_{W}=80.377\>{\rm GeV}$, $m_{t}=172.69\>{\rm GeV}$, $m_{e}=0.5109989\>{\rm MeV}$, $m_{\mu}=105.65837\>{\rm MeV}$, $m_{\tau}=1.77686\>{\rm GeV}$. The masses of the charged leptons are retained only when computing $Z_{e}$. In all other situations, we keep the top quark massive and treat the remaining five quarks and all charged leptons as massless.
We adopt the following values for various couplings: $\alpha(0)=1/137.035999$, $G_{\mu}=1.1663787\times 10^{-5}~{}{\rm GeV}^{-2}$, and $\Delta\alpha_{\rm had}^{(5)}(M_{Z})=0.02764$ Workman:2022ynf . The running strong coupling constant at the scale of the Higgs mass is set to $\alpha_{s}(M_{H})=0.115$ Bonciani:2015eua . With the aid of (8b), we derive $\alpha_{G_{\mu}}=1/132.168$. To calculate $\alpha(M_{Z})$, we must know the values of $\Delta\alpha_{q}(M_{Z})$ and $\Delta\alpha_{l}(M_{Z})$. As mentioned before, the nonperturbative parameter $\Delta\alpha_{q}(M_{Z})$ can be interchangeably used with the experimentally determined $\Delta\alpha_{\rm had}^{(5)}(M_{Z})$. $\Delta\alpha_{\ell}(M_{Z})$ can be calculated in perturbation theory, whose value is determined through four-loop accuracy Kuhn:1998ze ; Eidelman:1995ny ; Steinhauser:1998rq ; Sturm:2013uka . We adopt the four-loop result $\Delta\alpha_{\ell}\left(M_{Z}\right)\approx 0.0314979$ Sturm:2013uka . Consequently, by employing (8a), we obtain $\alpha(M_{Z})=1/128.932$.
At present the measured full width of the Higgs boson is subject to large uncertainty, $\Gamma_{H}=3.2^{+2.4}_{-1.7}$ MeV Workman:2022ynf . When predicting the branching fraction of $H\to Z^{0}\gamma$, we choose to use much more precise, theoretically predicted value of Higgs full width, $\Gamma_{H}=4.07^{+4.0\%}_{-3.9\%}$ MeV Workman:2022ynf ; LHCHiggsCrossSectionWorkingGroup:2016ypw , in order to reduce the theoretical error.
@@
$\alpha(0)$ scheme $\alpha^{3}(0)$ $5.921$ $0.313$ $0.019$ $6.252$ $1.54\pm 0.06$
$\alpha(M_{Z})$ scheme $\alpha(0)\alpha^{2}(M_{Z})$ $6.689$ $-0.464$ $0.021$ $6.245$ $1.53\pm 0.06$
$G_{\mu}$ scheme $\alpha(0)\alpha^{2}_{G_{\mu}}$ $6.365$ $-0.048$ $0.020$ $6.337$ $1.56\pm 0.06$
Democratic scheme $\alpha(0)\alpha(M_{Z})\alpha_{G_{\mu}}$ $6.525$ $-0.249$ $0.021$ $6.297$ $1.55\pm 0.06$
@@
Table 1: Predicted values of the partial width (in units of keV) for $H\to Z^{0}\gamma$ in various $\alpha$ schemes, at various levels of perturbative accuracy. $\Gamma^{\rm LO}$ represents the LO prediction to the decay width. The contributions from the NLO electroweak and QCD corrections are denoted by $\Gamma^{\mathcal{O}(\alpha)}$ and $\Gamma^{\mathcal{O}(\alpha_{s})}$, respectively. We have taken the strong coupling constant $\alpha_{s}(M_{H})=0.115$. $\Gamma^{\rm Sum}$ is obtained by summing the LO contribution together with $\mathcal{O}(\alpha)$ and $\mathcal{O}(\alpha_{s})$ corrections, and the rightmost column is the branching fraction obtained via dividing $\Gamma^{\rm Sum}$ by the predicted full Higgs width, $\Gamma_{H}=4.07^{+4.0\%}_{-3.9\%}$ MeV Workman:2022ynf ; LHCHiggsCrossSectionWorkingGroup:2016ypw .
@@
In Appendix A we present the numerical results of the form factor $\widetilde{T}_{5}$ at NLO accuracy in different $\alpha$ schemes. Plugging into (3), we are able to present the predictions for the partial width in various $\alpha$ schemes. It may look natural to take the QED coupling constant at the vertex attached to the outgoing photon as $\alpha(0)$. For the electromagnetic coupling constant that appears elsewhere, one has some freedom to choose different $\alpha$ scheme. In Table 1 we enumerate our predictions for $\Gamma(H\to Z^{0}\gamma)$ (in units of keV) based on four different $\alpha$ schemes at various levels of perturbative accuracy. For reader’s convenience, in Table 1 we also juxtapose the contribution of the NLO QCD correction to the partial width, which has been available long while ago Spira:1991tj ; Gehrmann:2015dua ; Bonciani:2015eua .
@@
As can be seen from Table 1, we reconfirm that the NLO QCD correction is positive and tiny, which constitutes only $0.3\%$ of the LO prediction of the partial width Bonciani:2015eua . We also notice that, the NLO electroweak corrections is positive in the $\alpha(0)$ scheme, while negative in other three schemes. In addition, the ${\cal O}(\alpha)$ corrections appear to be sizable in magnitude in both the $\alpha(0)$ and $\alpha(M_{Z})$ schemes, which may approach approximately ${}^{+5\%}_{-7\%}$ of the LO prediction. In contrast, the $G_{\mu}$ scheme as well as the Democratic scheme yield a relatively modest ${\cal O}(\alpha)$ correction. Note the LO predictions of the partial width from four different $\alpha$ scheme are scattered in a wide range, from $5.921$ keV to $6.689$ keV. Including the error of the Higgs full width, we find that the LO SM prediction for the branching fraction lies in the range $(1.40-1.71)\times 10^{-3}$. Interestingly, after including the NLO electroweak correction, the scheme dependence becomes substantially reduced, with the relative uncertainties among different schemes less than $2\%$. Taking the predictions from various $\alpha$ schemes as an estimation of the theoretical uncertainty, we then obtain the most accurate SM prediction to be ${\cal B}[H\to Z^{0}\gamma]=(1.55\pm 0.06)\times 10^{-3}$. Note the error in the branching fraction mainly stem from the uncertainty associated with the full width of the Higgs boson  [^6] We can estimate other potential sources of uncertainty. The contribution from the top quark loop constitutes approximately $-10\%$ of the LO decay width. Taking into account that the Yukawa coupling strength of $Hb\bar{b}$ is suppressed with respect to $Ht\bar{t}$ by a factor of $m_{b}/m_{t}$, we estimate that retaining the bottom quark mass may introduce a relative error of several per mille. Furthermore, uncertainties about the $t$, $W$ masses may also introduce an uncertainty of several per mille in the partial width. All in all, these sorts of uncertainties are of the same order of magnitude as the NLO QCD correction, which are significantly smaller than the uncertainty stemming from four different $\alpha$ schemes..
@@
It is evident that the state-of-the-art SM prediction for the branching fraction is significantly lower than the measured value, ${\cal B}_{\rm exp}[H\to Z^{0}\gamma]=(3.4\pm 1.1)\times 10^{-3}$ ATLAS:2023yqk . It may be too early to claim that some sort of the BSM physics must be invoked to resolve this discrepancy. More accurate measurements from HL-LHC, and the prospective Higgs factories such as CEPC and FCC-ee, will play a crucial role to ultimately clear the smoke.